OpenStudy (anonymous):

how do i write y=x^2+8x+15 in vertex form? please explain

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" or sometimes called a "perfect square" is?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok... then let's start by grouping \(\bf y=x^2+8x+15\implies y=(x^2+8x)+15\implies y=(x^2+8x+{\color{red}{ \square }}^2)+15\) do you know what number we're missing to get a perfect square trinomial in the group?

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

well... recall that 8 = 2 * 2 * 2 => 2 * 4 thus \(\bf y=x^2+8x+15\implies y=(x^2+8x)+15\implies y=(x^2+8x+{\color{red}{ \square }}^2)+15 \\ \quad \\ y=(x^2+2(4)(x)+{\color{red}{ \square }}^2)+15\) what about now?

OpenStudy (anonymous):

4?

OpenStudy (jdoe0001):

yeap thus now we know the 2nd term is "4" since the middle term of the perfect square is 2 * the other terms so keep in mind that all we're doing is borrowing from our friend zero, 0, so we'll ADD \(4^2\) BUT we also have to SUBTRACT \(4^2\) thus \(\bf y=x^2+8x+15\implies y=(x^2+8x)+15\implies y=(x^2+8x+{\color{red}{ \square }}^2)+15 \\ \quad \\ y=(x^2+2(4)(x)+{\color{red}{ 4 }}^2)+15-{\color{red}{ 4 }}^2 \\ \quad \\ y=(x+4)^2+15-16\implies y=(x+4)^2-1\impliedby vertex\ form\)

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw